# A Survey on Employing Large Language Models for Text-to-SQL Tasks

### v1

@@ 4.4. Reasoning
Similar to the application of LLM in other scenarios, The application of LLM in the Text2SQL field can adopt a multi-step reasoning pattern, generating answers that gradually approach the correct SQL step by step. The main multi-step reasoning methods include Chain-of-Thought (CoT)(Wei et al., 2022b), Least-to-Most(Zhou et al., 2022), Self-Consistency(Wang et al., 2022c) and Self-Correction.
@@ 4.4.1. Chain of Thought (CoT)
Chain-of-Thought(CoT) guides LLM to think step by step and generate answers by adding statements such as ”Let’s think step by step” to the prompt words, allowing LLM to execute multiple steps in sequence and generate the final result in the last step. A lot of work (Zhang et al., 2023a; Li et al., 2024a; Hu et al., 2023; Wang et al., 2023a; Pourreza and Rafiei, 2024; Liu and Tan, 2023; Wang et al., 2023b; Xie et al., 2024) adopted the CoT method to help LLM generate SQL. ACT-SQL (Zhang et al., 2023a) utilizes LLM’s few-shot method to automatically extend ordinary examples into CoT pattern examples. Through multiple rounds of LLM interaction, the SQL generated from the previous interaction will be executed and affect the in-memory database(Hu et al., 2023). The next interaction will use the current in-memory database status as prompt words, allowing LLM to decide the next database operation. DBCopilot (Wang et al., 2023a) guides LLM to select appropriate tables and columns before writing SQL in the example, allowing LLM to generate two-step CoT answers. In (Liu and Tan, 2023), three types of CoT solutions were proposed for LLM to solve the Text2SQL problem, namely clause-by-clause generation, selecting tables and columns before generation, and refining after completion.
@@ 5.2. Selection of Pretrained Models
Numerous open-source language models are available, such as (Touvron et al., 2023b; Roziere et al., 2023; Li et al., 2023a; Nijkamp et al., 2022; Guo et al., 2024; Anil et al., 2023). These models vary in terms of parameter scales, pre-training corpora, and pre-training processes, and all can be fine-tuned. For example, the PaLM-2(Anil et al., 2023) model was fine-tuned for the text2sql task as detailed in (Sun et al., 2023), and it achieved the best results on the Spider dataset at that time. The code-llama(Roziere et al., 2023) model was fine-tuned in (Wang et al., 2023b) and was used in the Agents system, which was designed for the text2sql task. However, when selecting a model for fine-tuning, it is crucial to consider a variety of factors, including the pre-trained model’s capabilities, the volume of fine-tuning data, the model parameters’ scale, and computational capabilities. As mentioned in (Chung et al., 2024), the advantages of fine-tuning may be attributed to its capacity to more effectively utilize the knowledge gained during pre-training. Consequently, selecting a language model with a more extensive pre-training corpus could be advantageous. (Dong et al., 2023a) indicates that pre-trained models with larger parameter scales perform better when given the same amount of training data. Research in (Zhang et al., 2024) found that coding-specific models are more effective in Text-to-SQL tasks.

### v2

@@ 4.4. Reasoning
Similar to the application of LLM in other scenarios, The application of LLM in the Text2SQL field can adopt a multi-step reasoning pattern, generating answers that gradually approach the correct SQL step by step. The main multi-step reasoning methods include Chain-of-Thought (CoT)(Wei et al., 2022b), Least-to-Most(Zhou et al., 2022), Self-Consistency(Wang et al., 2022b) and Self-Correction.
@@ 4.4.1. Chain of Thought (CoT)
Chain-of-Thought(CoT) guides LLM to think step by step and generate answers by adding statements such as ”Let’s think step by step” to the prompt words, allowing LLM to execute multiple steps in sequence and generate the final result in the last step. A lot of work (Zhang et al., 2023; Li et al., 2024a; Hu et al., 2023; Wang et al., 2023a; Pourreza and Rafiei, 2024; Liu and Tan, 2023; Wang et al., 2023b; Xie et al., 2024) adopted the CoT method to help LLM generate SQL. ACT-SQL (Zhang et al., 2023) utilizes LLM’s few-shot method to automatically extend ordinary examples into CoT pattern examples. Through multiple rounds of LLM interaction, the SQL generated from the previous interaction will be executed and affect the in-memory database(Hu et al., 2023). The next interaction will use the current in-memory database status as prompt words, allowing LLM to decide the next database operation. DBCopilot (Wang et al., 2023a) guides LLM to select appropriate tables and columns before writing SQL in the example, allowing LLM to generate two-step CoT answers. In (Liu and Tan, 2023), three types of CoT solutions were proposed for LLM to solve the Text2SQL problem, namely clause-by-clause generation, selecting tables and columns before generation, and refining after completion.
@@ 5.2. Selection of Pretrained Models
Numerous open-source language models are available, such as (Touvron et al., 2023; Roziere et al., 2023; Li et al., 2023a; Nijkamp et al., 2022; Guo et al., 2024; Anil et al., 2023). These models vary in terms of parameter scales, pre-training corpora, and pre-training processes, and all can be fine-tuned. For example, the PaLM-2(Anil et al., 2023) model was fine-tuned for the text2sql task as detailed in (Sun et al., 2023), and it achieved the best results on the Spider dataset at that time. The code-llama(Roziere et al., 2023) model was fine-tuned in (Wang et al., 2023b) and was used in the Agents system, which was designed for the text2sql task. However, when selecting a model for fine-tuning, it is crucial to consider a variety of factors, including the pre-trained model’s capabilities, the volume of fine-tuning data, the model parameters’ scale, and computational capabilities. As mentioned in (Chung et al., 2024), the advantages of fine-tuning may be attributed to its capacity to more effectively utilize the knowledge gained during pre-training. Consequently, selecting a language model with a more extensive pre-training corpus could be advantageous. (Dong et al., 2023a) indicates that pre-trained models with larger parameter scales perform better when given the same amount of training data. Research in (Zhang et al., 2024) found that coding-specific models are more effective in Text-to-SQL tasks.